# A1: Asynchronous Test-Time Scaling via Conformal Prediction

### v1

@@ Appendix A Appendix
A.1 Related Work
@@ A.2 Experimental Setup
Our evaluation covers four reasoning benchmarks. We use 100 randomly sampled problems from MATH (Hendrycks et al., 2021) (denoted as MATH100) for grade-school arithmetic word problems, AIME24 (Art of Problem Solving, ) and AIME25 (OpenCompass, ) for high-school competition-level mathematics, and the first 50 problems from AMC23 (math-ai, ) for the American Mathematics Competitions. These datasets require multi-step reasoning and are particularly suitable for testing the effectiveness of asynchronous sampling with rejection. To ensure consistency, we set a token budget of 8192 across all settings and adopt deterministic decoding with temperature set to zero.
@@
A.3 Controlling the rejection rate in multi-turn interactions

### v2

@@ Appendix A Appendix
A.1 The Use of Large Language Models
In accordance with the ICLR policy on the use of large language models, we hereby declarethat LLMs were employed solely to assist in improving the grammar and enhancing the expressionof this paper. The original research idea, methodological development, and overall structure andcontent of the manuscript were entirely conceived and written by the authors. At no stage was theuse of LLMs extended to the generation of core intellectual content, and we afirm that there hasbeen no misuse of LLMs in the preparation of this work.
A.2 Related Work
@@ A.3 Experimental Setup
Our evaluation covers four reasoning benchmarks. We use 100 randomly sampled problems from MATH (Hendrycks et al., 2021) (denoted as MATH100) for grade-school arithmetic word problems, AIME24 (Art of Problem Solving, ) and AIME25 (OpenCompass, ) for high-school competition-level mathematics, and the first 50 problems from AMC23 (math-ai, ) for the American Mathematics Competitions. These datasets require multi-step reasoning and are particularly suitable for testing the effectiveness of asynchronous sampling with rejection. To ensure consistency, we set a token budget of 8192 across all settings and adopt deterministic decoding with temperature set to zero. We set the maximum number of turns to 10.
@@
Figure 9: Accuracy improvement with increasing sample size on AIME 2025.
A.4 Large-scale asynchronous test-scaling
In this section, we set the maximum number of test-time scaling turns of the model to 20, and then gradually increase the number of samples per turn up to 128. We conduct comparative experiments under editing coverage, conditional coverage, as well as under the standard rejection sampling and our designed rejection sampling settings. We set the draft model to DeepSeek-R1-Distill-Qwen-1.5B, and the target model to DeepSeek-R1-Distill-Llama-70B.
According to Figure 9, we can observe that as the number of samples increases, the performance of our model gradually improves. With 8 samples, our model reaches the performance of o3-mini(low); with 64 samples, it reaches the performance of o3-mini(medium); and with 128 samples, it reaches the performance of o3-mini(high), which are closed-source reasoning models. Due to limited computational resources, we did not conduct experiments with larger-scale sampling, which results in still falling short of GPT-5 performance.
Continue Sampling.
In our experimental setup, we did not strictly adhere to the standard rejection sampling procedure (i.e., discarding the samples generated by the draft model and resampling with the target model) when performing scaling. Instead, under the continue sampling setting, if a sample produced by the draft model is included in the prediction set of the current turn, the target model subsequently continues the sampling in the following turn conditioned on this sampled result. According to Figure 9, both our conditional coverage and editing coverage adopt the continue-sampling scheme. The conditional coverage demonstrates relatively high sampling efficiency, reaching the level of o3-mini-medium under the 128-sample setting.
Resampling.
We also conducted experiments that scale the number of samples under the standard rejection sampling setting. In this setting, at each scaling turn, if the draft model’s sample is included in the prediction set for the current turn, then within the same turn the target model draws a prediction set whose size matches that of the current prediction set. As indicated by the red curve in Figure 9, this scheme exhibits substantially higher sampling efficiency than the alternatives; however, for the same nominal number of samples it consumes more tokens (since part of the draft model’s tokens are discarded). Under the 128-sample setting, it achieves performance comparable to o3-mini-high.
(a) The change in final accuracy as $\alpha$ increases.
(b) The increase in total time overhead as $\alpha$ grows.
Figure 10: Ablation study on the hyperparameter $\alpha$ using the DeepSeek-Qwen -1.5B and QwQ-32B models on the AIME2025 dataset.
A.5 Ablation Study
In this section, we present the ablation experiments on the hyperparameter $\alpha$. As shown in Figure 10, as the hyperparameter $\alpha$ increases, the overall accuracy and time overhead of the system both rise. This reflects that the intervention of the target model increases both accuracy and time overhead. However, we can find a balance between accuracy and time overhead, such as when $\alpha$ is 0.2, where the condition coverage setting achieves 60% accuracy with relatively low time overhead. When $\alpha$ is 0.4, both the marginal coverage and condition coverage settings reach 60% accuracy, making it a more robust hyperparameter setting.
@@ A.5 Ablation Study
A.6 Controlling the rejection rate in multi-turn interactions